# Integral modelling and Reinforcement Learning control of 3D liquid metal coating on a moving substrate

### v1

@@ III.3 The long wave formulation
where $\Upsilon=\mbox{Ha}\,\hat{b}$ and the coefficients $c_{1},c_{2},c_{3},c_{4},c_{5},c_{6}$ depend on $\hat{h},\hat{q}_{x}$, $\hat{q}_{z}$, Ha, and $\hat{b}$. By solving a linear system of equations, which includes the non-slip condition at the substrate (20a), the tangential stress balance at the free surface (20d), and the flow rate definitions (3), the following expressions for these coefficients are derived:
$\begin{split}c_{1}=\frac{-\hat{h}\Upsilon^{2}+\Xi e^{\hat{h}\Upsilon}(\hat{h}\Upsilon-1)-\hat{q}_{x}\Upsilon^{2}+\Xi}{\Upsilon\left(\hat{h}\Upsilon+e^{2\hat{h}\Upsilon}(\hat{h}\Upsilon-1)+1\right)}\,,\end{split}$ (26a)
$\begin{split}c_{2}=\frac{e^{2\hat{h}\Upsilon}\left(\Xi-\Upsilon^{2}(\hat{h}+\hat{q}_{x})\right)-\Xi e^{\hat{h}\Upsilon}(\hat{h}\Upsilon+1)}{\Upsilon\left(\hat{h}\Upsilon+e^{2\hat{h}\Upsilon}(\hat{h}\Upsilon-1)+1\right)}\,,\end{split}$ (26b)
$\begin{split}c_{3}=\frac{\Upsilon\left(\cosh(\hat{h}\Upsilon)\left(\hat{q}_{x}\Upsilon^{2}-\Xi\right)+\Upsilon\sinh(\hat{h}\Upsilon)+\Xi\right)}{\hat{h}\Upsilon\cosh(\hat{h}\Upsilon)-\sinh(\hat{h}\Upsilon)}\,,\end{split}$ (26c)
$\begin{split}c_{4}=\frac{\Psi e^{\hat{h}\Upsilon}(\hat{h}\Upsilon-1)+\Psi-\hat{q}_{z}\Upsilon^{2}}{\Upsilon\left(\hat{h}\Upsilon+e^{2\hat{h}\Upsilon}(\hat{h}\Upsilon-1)+1\right)}\,,\end{split}$ (26d)
$\begin{split}c_{5}=\frac{e^{2\hat{h}\Upsilon}\left(\Psi-\hat{q}_{z}\Upsilon^{2}\right)-\Psi e^{\hat{h}\Upsilon}(\hat{h}\Upsilon+1)}{\Upsilon\left(\hat{h}\Upsilon+e^{2\hat{h}\Upsilon}(\hat{h}\Upsilon-1)+1\right)}\,,\end{split}$ (26e)
@@
$\begin{split}&F_{12}=\Big{[}e^{2\hat{h}\Upsilon}(\sinh(2\hat{h}\Upsilon)(\Xi^{2}(\hat{h}^{2}\Upsilon^{2}+2)\\ &+\Upsilon^{4}(\hat{h}-\hat{q}_{x})(\hat{h}+3\hat{q}_{x})+2\Upsilon^{2}\Xi(\hat{h}+\hat{q}_{x}))\\ &+2\Upsilon^{3}(\hat{h}^{3}(-\Xi^{2})+\hat{h}\Upsilon^{2}(\hat{h}^{2}+2\hat{h}\hat{q}_{x}+2\hat{q}_{x}^{2})\\ &-2\hat{h}\Xi(\hat{h}+\hat{q}_{x})+\hat{h}+2\hat{q}_{x})\\ &+2\Upsilon\cosh(2\hat{h}\Upsilon)(\hat{h}\hat{q}_{x}^{2}\Upsilon^{4}-\Upsilon^{2}(\hat{h}+2\hat{q}_{x})-2\hat{h}\Xi^{2})\\ &-4\Xi\sinh(\hat{h}\Upsilon)(\Upsilon^{2}(\hat{h}(\hat{h}\Upsilon^{2}(\hat{h}+\hat{q}_{x})-\hat{h}\Xi+1)\\ &+\hat{q}_{x})+\Xi)+4\hat{h}\Upsilon\Xi\cosh(\hat{h}\Upsilon)(\Upsilon^{2}(\hat{h}+\hat{q}_{x})\\ &+\Xi))\Big{]}/\Big{[}\Upsilon^{3}(\hat{h}\Upsilon+e^{2\hat{h}\Upsilon}(\hat{h}\Upsilon-1)+1)^{2}\Big{]}\,,\end{split}$ (27a)
$\begin{split}&F_{13}=\Big{[}e^{2\hat{h}\Upsilon}(-2\sinh(\hat{h}\Upsilon)(\Psi\Upsilon^{2}(\hat{h}+\hat{q}_{x})(\hat{h}^{2}\Upsilon^{2}+1)\\ &+\Xi(-2\hat{h}^{2}\Psi\Upsilon^{2}+\hat{h}^{2}\hat{q}_{z}\Upsilon^{4}+2\Psi+\hat{q}_{z}\Upsilon^{2}))\\ &+\sinh(2\hat{h}\Upsilon)(\Upsilon^{2}\Xi(\hat{h}^{2}\Psi+\hat{q}_{z})+\Psi\Upsilon^{2}(\hat{h}+\hat{q}_{x})\\ &+\hat{q}_{z}\Upsilon^{4}(\hat{h}-3\hat{q}_{x})+2\Psi\Xi)-2\Upsilon^{3}(\hat{h}(\hat{h}^{2}\Psi\Xi+\hat{h}\Psi\\ &-\hat{h}\hat{q}_{z}\Upsilon^{2}+\Psi\hat{q}_{x}-2\hat{q}_{x}\hat{q}_{z}\Upsilon^{2}+\hat{q}_{z}\Xi)-\hat{q}_{z})\\ &+2\hat{h}\Upsilon\cosh(\hat{h}\Upsilon)(\Psi\Upsilon^{2}(\hat{h}+\hat{q}_{x})+\Xi(2\Psi+\hat{q}_{z}\Upsilon^{2}))\\ &+2\Upsilon\cosh(2\hat{h}\Upsilon)(-2\hat{h}\Psi\Xi+\hat{h}\hat{q}_{x}\hat{q}_{z}\Upsilon^{4}-\hat{q}_{z}\Upsilon^{2}))\Big{]}\\ &/\Big{[}\Upsilon^{3}(\hat{h}\Upsilon+e^{2\hat{h}\Upsilon}(\hat{h}\Upsilon-1)+1)^{2}\Big{]}\,,\end{split}$ (27b)
@@
with $F_{22}=F_{13}$ and for the wall shear stress term $\hat{\bm{\tau}}_{w}=(\hat{\tau}_{w,x},\hat{\tau}_{w,z})^{T}$, reading:
@@
Equations (22) with the closure relations (27) and (28) define the 3D IBL model with gas jets and electromagnetic actuators. Considering the conditions relevant to this analysis (see Table 1), we neglect Maxwell stresses as their normal $\Lambda\hat{h}/2$ and tangential $\hat{\bm{\tau}}_{M}=(\Lambda\hat{b}^{2}\partial_{x}\hat{h},\Lambda\hat{b}^{2}\partial_{z}\hat{h})$ components appear at $O(10^{-1})$ in the equations. Despite $\Lambda=9$, and given that $\hat{b}\in[0,1]$, $\hat{h}\ll 1$, with derivatives satisfying $\partial_{x}\hat{h}=O(1)$ and $\partial_{z}\hat{h}=O(1)$, we can confidently neglect these terms.
@@ IV Modelling control actuators
This section outlines the modelling of the gas jet and magnetic forces acting on the liquid film, which are included in the reduced-dimensional model equations (22) and the closure relations (27) and (28) as pressure $p_{g}$ and shear stress $\bm{\tau}_{g}$ distributions at the free surface and as a wall-normal magnetic field $\hat{b}$. Pressure $p_{g}$ and shear stress $\bm{\tau}_{g}$ are modelled using experimental correlations for circular impinging gas jets on a dry substrate (Subsection IV.1). The magnetic field $b$ is modelled as a Gaussian distribution, which approximates the magnetic field of a cylindrical electromagnetic actuator (Subsection IV.2).
IV.1 Modelling gas jets actuators
Figure 2: Schematic of the impinging gas jet emerging from a circular nozzle with diameter $d$ and centreline velocity $U_{j}$, positioned at a distance $H$ from the substrate.
The distribution of pressure $\hat{p}_{\rm g}$ and shear stresses $\hat{\bm{\tau}}_{\rm g}=(\hat{\tau}_{{\rm g},x},\hat{\tau}_{{\rm g},z})^{T}$ on the free surface is modelled using experimental correlations for circular gas jets impinging on a flat dry substrate [40]. Figure 2 provides a schematic of the gas jet impingement configuration and its relevant parameters. In this study, we consider nozzles with a diameter of $d=1.5\,\rm{mm}$ positioned at a distance $H=1.4\,\rm{cm}$ from the substrate. The assumption of one-way coupling is justified by the large ratio $H\gg h$, as demonstrated in Lacanette et al. [41] and Gosset et al. [20] for the average film distribution. Following Mendez et al. [17], this assumption is also regarded appropriate for the dynamic case to demonstrate the feasibility of electromagnetic control.
Figure 2 also shows the location of the radial reference system $\{\mathbf{O},r\}$, positioned on the substrate along the centreline of the jet.
The value of the stagnation pressure $p_{s}$ (at $r=0$) and the maximum shear stress $\tau_{0m}$ (at $r_{1}=0.14\,H$) are given by:
$p_{s}=50\frac{\rho_{g}U_{j}^{2}}{2}\Big{(}\frac{d}{H}\Big{)}^{2},\quad\tau_{0m}=0.16\rho_{g}U_{j}^{2}\Big{(}\frac{d}{H}\Big{)}^{2},$ (29)
where $U_{j}$ is the centreline nozzle’s exit velocity.
@@ IV.1 Modelling gas jets actuators
The experimental correlations for the pressure $p_{g}(r)$ and shear stress $\tau_{g}(r)$ distributions read:
$p_{g}=p_{s}e^{(-114\lambda^{2})},$ (30a)
$\begin{split}\tau_{g,{\rm mod}}=\max\Big{(}0,&\tau_{0m}(0.18\frac{(1-e^{-114\lambda^{2}})}{\lambda}+\\ &-9.43\lambda e^{-114\lambda^{2}}-\lambda^{2})\Big{)},\end{split}$ (30b)
where $\lambda=r/H$ and $\tau_{g,{\rm mod}}$ is a modified form of the original shear stress distribution proposed by Beltaos and Rajaratnam [40], including an additional quadratic term in $\lambda$ and a clipping function between zero and $\tau_{0m}$.
Figure 3: Pressure (blue continuous line with circles) and shear stress (green dash-dotted line with triangles) distributions at the strip level in the radial direction $\lambda=r/H$ for a circular 3D jet with the modified shear stress distribution (orange dashed line with squares) and a highlight of the impingement and the wall jet regions.
Figure 3 shows the normalised pressure (solid blue line with circles) and shear stress (orange dashed line with squares) distributions, compared to the shear stress distribution proposed by Beltaos and Rajaratnam [40] (dash-dotted line with triangles) and the shear stress measurements from Beltaos and Rajaratnam [40] (black diamonds) and Bradshaw and Love [42] (red stars). The stress distributions are more pronounced in the impingement region (grey area) ($\lambda\leq 2.2$), and gradually decrease to zero in the wall jet region ($\lambda>2.2$). Our modified shear stress distribution, given by (30b), more accurately predicts the experimental data than Bradshaw’s correlation, which deviates from observed values outside the impingement region.
@@ IV.2 Modelling electromagnetic actuators
The electromagnetic actuators are modelled as cylindrical solenoids perpendicular to the substrate with length $L_{s}$, radius $R_{s}$, and metal core with magnetic permeability $\mu_{m}$. Figure 4 shows a solenoid schematic containing $n$ coils and carrying a time-varying current $I(t)$. In a cylindrical reference frame $\{\mathbf{O};r_{s},z_{s}\}$ centred at the midpoint of the solenoid, the magnetic field is given by complex analytical expressions derived by Hampton et al. [43] (see Appendix B for more details).
Figure 4: Scheme of the solenoid of length $L_{s}$, radius $R_{s}$ with a current $I$ with the cylindrical reference frame centred in the midpoint.
Assuming that the liquid film is positioned at a distance $z\ll R_{s}-L_{s}/2$ and considering a region near the centerline ($r_{s}\to 0$), the radial component of the magnetic field can be neglected and its axial component ($\hat{b}(\hat{x},\hat{z},\hat{t})$ in the reference frame of the liquid film) can be approximated as a 2D Gaussian in the $(x,z)$ plane, with a time-dependent modulation described by the control function $\hat{b}_{t}(\hat{t})$. The approximated magnetic field reads:
$\hat{b}(\hat{x},\hat{z},\hat{t})=\hat{b}_{t}(\hat{t})\,\exp\left(\frac{-(\hat{x}-\hat{x}_{0})^{2}}{2\hat{\gamma}^{2}}+\frac{-(\hat{z}-\hat{z}_{0})^{2}}{2\hat{\gamma}^{2}}\right),$ (31)
where ($x_{0},z_{0}$) denotes the position of the Gaussian’s centerline and $\hat{\gamma}$ is the nondimensional standard deviation. The dimensional value of $\gamma$ defines the spatial scale at which the electromagnet affects the liquid film ($bx_{\rm ref}=\gamma x_{\rm ref}$). This value is important for calculating the characteristic timescale of the magnetic effect according to (12).
The maximum frequency $\hat{f}_{b,\max}$ of $\hat{b}_{t}(\hat{t})$ is constrained to avoid modeling the skin effect. This effect consists in an increase of the liquid film’s electrical resistivity due to electric currents concentrating in a thin region near the free surface referred to as the skin thickness $h_{S}$, reading [44, Section 11.4]:
$h_{S}=\sqrt{\frac{t_{\rm ref}}{\pi\hat{f}_{b,\max}\mu_{M}\sigma_{M}}}.$ (32)
This thickness represents the depth within the liquid film at which the amplitude of an electromagnetic wave decreases to $1/e$ (approximately 37%) of its initial value at the surface. To ensure that the skin effect can be neglected, we require that $h_{S}$ remains larger than the liquid film thickness $h$, leading to an upper bound on for $\hat{f}_{b,\max}$, given by:
$\hat{f}_{b,{\rm max}}<\frac{t_{\rm ref}}{\pi\mu_{M}\sigma_{M}h_{\rm ref}^{2}\hat{h}^{2}}\approx\frac{658}{\hat{h}^{2}}\,.$ (33)
To minimize the error between the approximated form (31) and the analytical solution [43], we derived a relation between the nondimensional amplitude $\hat{b}_{t}$, the standard deviations $\hat{\gamma}$, and the geometrical and physical characteristics of the solenoid (detailed in Appendix B), which read:
$\hat{b}_{t}=\frac{\mu_{m}I(t)nL_{s}}{4R_{s}b_{\rm ref}\mbox{Ca}^{1/3}},\quad\quad\hat{\gamma}=\frac{R_{s}}{2^{3/4}\sqrt{3}\sqrt[4]{17}h_{\rm ref}}.$ (34)
These relations can help design an experimental setup: given the shape and intensity of the magnetic field $b(x,z,t)$ used in our simulations, they allow for defining the electromagnet geometry, the number of coils, and the time-dependent current profile in the experimental setup.
@@ V Numerical Methods
The IBL model in (22) was implemented in numerical software using the Fourier pseudo-spectral method for spatial discretization [45, 46]. This method transforms the physical quantities from a physical space to a Fourier wave number space, where the spatial derivatives are calculated analytically. The liquid film height $\hat{h}$ and the flow rates $\hat{q}_{\hat{x}}$ and $\hat{q}_{\hat{z}}$ are approximated as linear combinations of $N\times M$ 2D plane waves, where $N,M\in\{2n\mid n\in\mathbb{N}^{+}\}$, expressed as:
@@
Substituting (35) into (22), calculating the spatial derivative by differentiating (35) and subtracting the right-hand-side from the left-hand-side results gives the residual of the spectral approximation. This is then projected onto a basis of Dirac delta functions $\delta_{\hat{x}_{i},\hat{z}_{j}}$, defined on an equispaced grid (known as collocation points) where $\hat{x}_{i}=\frac{2\pi}{N}i$ and $\hat{z}_{j}=\frac{2\pi}{M}j$, with $i,j\in\mathbb{N}$ such that $i\leq N$ and $j\leq M$. Using the properties of the Dirac delta function, the inner product with the residual yields a discrete inverse Fourier transform (IDFT), which provides the values of the state variables at the grid points:
@@
Given the values of $\hat{h}$, $\hat{q}_{x}$, and $\hat{q}_{z}$ at the grid points for a specific time $\hat{t}$, the nonlinear terms are first computed in physical space and then transformed into Fourier wave number space using the Fast Fourier Transform (FFT). A 2D Finite Impulse Response (FIR) filter is applied to reduce aliasing errors, with a normalized cutoff wave number set to $0.8$ times the Nyquist frequency in both spatial directions. The filter is designed using a Hamming window, with the number of filter coefficients set to half the length of the input signal in each direction. The 2D filter is obtained by independently applying the dot product of the 1D filters along the $x$ and $z$ directions. Finally, the derivatives are transformed back into physical space using the Inverse Discrete Fourier Transform (IDFT), and time integration is carried out using Euler’s method.

### v2

@@ III.3 The long wave formulation
where $\Upsilon=\mbox{Ha}\,\hat{b}$ and the coefficients $c_{1},c_{2},c_{3},c_{4},c_{5},c_{6}$ depend on $\hat{h},\hat{q}_{x}$, $\hat{q}_{z}$, Ha, and $\hat{b}$. By solving a linear system of equations, which includes the non-slip condition at the substrate (25), the tangential stress balance at the free surface (26c), and the flow rate definitions (3), the following expressions for these coefficients are derived:
[…]
@@
[…]
@@
with $F_{22}=F_{13}$ and for the wall shear stress term $\hat{\boldsymbol{\tau}}_{w}=(\hat{\tau}_{w,x},\hat{\tau}_{w,z})^{T}$, reading:
@@
Equations (28) with the closure relations (33) and (34) define the 3D IBL model with gas jets and electromagnetic actuators.
@@ IV Modelling control actuators
This section outlines the modelling of the gas jet and magnetic actuators, which appear in the reduced-dimensional model equations (28) and in the closure relations (33) and (34) as pressure ($p_{g}$) and shear stress ($\boldsymbol{\tau}_{g}$) distributions at the free surface, and as a wall-normal magnetic field ($\hat{b}$). The pressure $p_{g}$ and shear stress $\boldsymbol{\tau}_{g}$ are modelled by adapting experimental correlations (Subsection IV.1), while the magnetic field $b$ is represented as an approximation of the field induced by an electromagnet (Subsection IV.2).
IV.1 Modelling gas jet actuators
The pressure $\hat{p}_{\rm g}$ and shear stress $\hat{\boldsymbol{\tau}}_{\rm g}=(\hat{\tau}_{{\rm g},x},\hat{\tau}_{{\rm g},z})^{T}$ distributions at the free surface are modelled using experimental correlations originally developed for circular gas jets impinging on a dry flat substrate [4]. Although the present study considers two-dimensional (2D) gas jets, these correlations are convenient for the demonstrational purposes of this work and facilitate a straightforward adaptation from round to planar jet configurations.
We consider a 2D gas jet actuator with an equivalent nozzle diameter of $d=1.5,\mathrm{mm}$, positioned at a distance $H=1.4,\mathrm{cm}$ from the substrate. Since $H$ is much larger than the characteristic film thickness, the gas flow can be assumed to be decoupled from the liquid film dynamics, as shown in Lacanette et al. [27] and Gosset et al. [19]. Under this assumption, $\hat{p}_{\rm g}$ and $\hat{\boldsymbol{\tau}}_{\rm g}$ depend solely on the jet parameters and geometry, and are independent of the film evolution.
@@ IV.1 Modelling gas jet actuators
The correlations from Beltaos and Rajaratnam [4] define the pressure and shear stress distributions as:
$p_{g}=p_{s},f_{p},\qquad\qquad\tau_{g}=\max\left(0,\tau_{0m}f_{\tau}\right),$ (35)
where $p_{s}$ and $\tau_{0m}$ are the stagnation pressure and maximum wall shear stress, respectively, and $f_{p}$ and $f_{\tau}$ are functions of the radial coordinate $r=x-x_{i}$, measured from the jet impingement point $x_{i}$.
The characteristic scales $p_{s}$ and $\tau_{0m}$ are given by:
$p_{s}=50\frac{\rho_{g}U_{j}^{2}}{2}\left(\frac{d}{H}\right)^{2},\;\tau_{0m}=0.16\rho_{g}U_{j}^{2}\left(\frac{d}{H}\right)^{2},$ (36)
where $U_{j}$ is the jet exit velocity and $\rho_{g}$ the gas density.
The nondimensional spatial distributions read:
$f_{p}$ $=\exp(-114\,\lambda^{2}),$ (37a)
$f_{\tau}$ $=0.18\,\frac{1-\exp(-114\,\lambda^{2})}{\lambda}$
$\quad-9.43\,\lambda\,\exp(-114\,\lambda^{2})-\lambda^{2},$ (37b)
where $\lambda=r/H$. Equation (37) represents a refined version of the original shear stress correlation, including an additional quadratic term for improved accuracy. To model 2D jets, the radial coordinate $r$ is replaced by the coordinate normal to the jet symmetry plane. In this transformation, the pressure distribution $p_{g}$ remains symmetric with respect to the stagnation line, while the shear stress $\tau_{g}$ becomes antisymmetric to reflect the change in flow direction across the impingement point, i.e.
$\tau_{g}(x)=\mathrm{sign}(x-x_{i})\,\tau_{0m}\,f_{\tau}(|x-x_{i}|).$ (38)
This antisymmetric formulation ensures a physically consistent adaptation of the circular jet correlation to a planar jet configuration.
@@ IV.2 Modelling electromagnetic actuators
The magnetic field $\hat{b}$, induced by the 2D electromagnetic actuators, is obtained by simplifying the analytical expression for a magnetic field induced by a cylindrical solenoid perpendicular to the substrate with length $L_{s}$, radius $R_{s}$, and a metal core with magnetic permeability $\mu_{m}$. In a cylindrical reference frame $\{\mathbf{O};r_{s},z_{s}\}$ centred at the midpoint of the solenoid, the magnetic field is given by complex analytical expressions derived by Hampton et al. [20] (see Appendix B for more details).
[…]
The 2D approximated magnetic field centred at $x_{0}$, reads:
$\hat{b}(\hat{x},\hat{t})=\hat{b}_{t}(\hat{t})\,\exp\left(\frac{-(\hat{x}-\hat{x}_{0})^{2}}{2\hat{\gamma}^{2}}\right),$ (39)
where $\hat{\gamma}$ is the nondimensional standard deviation. The dimensional value of $\gamma$ defines the spatial scale at which the electromagnet affects the liquid film ($bx_{\rm ref}=\gamma x_{\rm ref}$). This value is crucial for calculating the characteristic timescale of the magnetic effect, as defined in (15).
The control function $\hat{b}_{t}$ and the standard deviations $\hat{\gamma}$ are expressed as a function of the geometrical and physical characteristics of the solenoid (detailed in Appendix B), which read:
$\hat{b}_{t}=\frac{\mu_{m}I(t)nL_{s}}{4R_{s}b_{\rm ref}\mbox{Ca}^{1/3}},\quad\quad\hat{\gamma}=\frac{R_{s}}{2^{3/4}\sqrt{3}\sqrt[4]{17}h_{\rm ref}},$ (40)
where $n$ is the number of coils and $I(t)$ is time-varying current flowing through the system.
@@ V Numerical Methods
The IBL model in (28) was implemented in numerical software using the Fourier pseudo-spectral method for spatial discretisation [15, 9]. This method transforms physical quantities from physical space to Fourier wave-number space, in which spatial derivatives are computed via Fourier differentiation. The liquid film height $\hat{h}$ and the flow rates $\hat{q}_{\hat{x}}$ and $\hat{q}_{\hat{z}}$ are approximated as linear combinations of $N\times M$ 2D plane waves, where $N,M\in\{2n\mid n\in\mathbb{N}^{+}\}$, expressed as:
@@
Substituting (41) into (28), calculating the spatial derivative by differentiating (41) and subtracting the right-hand-side from the left-hand-side gives the residual of the spectral approximation. This is then projected onto a basis of Dirac delta functions $\delta_{\hat{x}_{i},\hat{z}_{j}}$, defined on an equispaced grid (collocation points) where $\hat{x}_{i}=\frac{2\pi}{N}i$ and $\hat{z}_{j}=\frac{2\pi}{M}j$, with $i,j\in\mathbb{N}$ such that $i\leq N$ and $j\leq M$. Using the properties of the Dirac delta function, the inner product with the residual yields a discrete inverse Fourier transform (IDFT), which provides the values of the state variables at the grid points:
@@
Given the values of $\hat{h}$, $\hat{q}_{x}$, and $\hat{q}_{z}$ at the grid points for a specific time $\hat{t}$, the nonlinear terms are first computed in physical space and then transformed into Fourier wave number space using the Fast Fourier Transform (FFT). A 2D Finite Impulse Response (FIR) filter is applied to reduce aliasing errors, with a normalised cutoff wave number set to $0.8$ times the Nyquist frequency in both spatial directions. The filter is designed using a Hamming window, with the number of filter coefficients set to half the length of the input signal in each direction. The 2D filter is obtained by independently applying the 1D filters along the $x$ and $z$ directions. Finally, the derivatives are transformed back into physical space using the Inverse Discrete Fourier Transform (IDFT), and time integration is carried out using Euler’s method.